### v2

# Client-Aided Secure Two-Party Computation of Dynamic Controllers

Abstract
In this paper, we propose a secure two-party computation protocol for dynamic controllers using a secret sharing scheme. The proposed protocol realizes outsourcing of controller computation to two servers, while controller parameters, states, inputs, and outputs are kept secret against the servers. Unlike previous encrypted controls in a single-server setting, the proposed method can operate a dynamic controller for an infinite time horizon without controller state decryption or input re-encryption. We show that the control performance achievable by the proposed protocol can be made arbitrarily close to that attained by the unencrypted controller. Furthermore, system-theoretic and cryptographic modifications of the protocol are presented to improve the communication complexity. The feasibility of the protocol is demonstrated through numerical examples of PID and observer-based controls.
{IEEEkeywords}
Cyber-physical systems, encrypted control, multi-party computation, privacy, security
1 Introduction
[…]
The need for protecting privacy and security in control systems has become critical due to growing cyber-physical systems in various industries. Encrypted control, which integrates automatic control with cryptographic technology, has emerged to prevent eavesdropping attacks and securely outsource controller computations for an untrusted third party [1, 2, 3]. The initial realization of encrypted control was achieved in a single-server setting by using homomorphic encryption [4]. Homomorphic encryption is an encryption scheme that allows arithmetic operations such as addition and multiplication on encrypted data [5, 6, 7, 8, 9]. In this framework, as illustrated in Fig. 11(a), the server receives encrypted measurements from the sensor, performs controller computation using encrypted parameters without decryption, and returns encrypted control actions to the actuator. Unlike other privacy mechanisms, such as differential privacy [10], encrypted control offers reliable security built on cryptography without a tradeoff between control performance and security level [3].
With existing homomorphic encryption schemes, arithmetic operations on encrypted data are usually performed on a finite range of integers. One of the main challenges in this context is to prevent overflow in recursive state updates when encrypting dynamic controllers (see Section 3). To address the overflow issue, fully homomorphic encryption with bootstrapping was employed in [11], and the periodic reset of controller states was considered in [12]. Recent approaches to this issue involve reformulating controller representations. For instance, one approach approximated a controller using a finite impulse response filter [13]. In addition, another approach reformulated a controller based on pole placement and coordinate transformation using re-encrypted control inputs [14]. The input re-encryption was also applied in [15] and [16], which represented controller states as historical input and output data.
(a) Encrypted control in a single-server setting.
(b) Encrypted control in a two-party computation setting.
Figure 1: System architectures of encrypted controls in single-server and two-party computation settings.
While these methods have advanced encrypted control in a single-server setting, exploring encrypted control in a multi-party computation setting presents potential advantages. Multi-party computation using secret sharing [17] is generally more computationally efficient and secure than homomorphic encryption in exchange for an increase in communication costs. Furthermore, multi-party computation offers the ability to evaluate various functionalities. This property motivates the implementation of encrypted control in a multi-party setting because homomorphic encryption typically requires substantial computational resources for computing mathematical operations beyond addition and multiplication.
Fig. 11(b) illustrates a typical system architecture of two-party encrypted control using secret sharing. Each controller receives a share of sensor measurements, securely performs controller computation (along with communication to the other controller if needed), and returns a share of control inputs to the actuator. In this framework, an efficient and secure implementation of static controllers was suggested [18]. This method was extended to polynomial control by combining secret sharing with homomorphic encryption [19]. The polynomial control was also realized based only on secret sharing in more than a three-party setting [20]. Furthermore, in a two-party setting, secret sharing with a garbled circuit was applied to implement a neural-network-based controller [21], and model predictive control using homomorphic encryption was considered [22].
In this paper, we aim to design an efficient protocol for securely implementing dynamic controllers. For this purpose, we follow the multi-party computation approach and perform bit truncation over secret sharing to prevent controller state overflow due to recursive multiplication. The main contribution of this study is to demonstrate the effectiveness of multi-party computation for encrypted dynamic control. Our contributions are summarized as follows:
  • We realize secure two-party computation of any linear time-invariant controllers without requiring state decryption or input re-encryption. This requirement of previous encrypted dynamic controls in a single-server setting [14, 15, 16] is undesirable because of increasing computation and communication efforts on a client. Compared to the previous schemes, the proposed protocol is simpler and significantly reduces the computation costs of a client.
  • We analyze the achievable control performance of the proposed protocol. This analysis clarifies that the performance can be arbitrarily close to that of the original controller by increasing the precision of the fixed-point numbers approximating the controller parameters and initial state. The result is examined by numerical examples with PID and observer-based controllers.
  • The proposed protocol guarantees statistical security as long as the two parties do not collude with each other. It is stronger than computational security satisfied by conventional encrypted controls using homomorphic encryption.
One of the main drawbacks of the proposed protocol is a high communication cost when implementing large-scale controllers. To attenuate communication efforts, we also provide modifications to the proposed protocol based on system-theoretic and cryptographic techniques.
The remainder of this paper is organized as follows. Section 2 describes a problem setting. Section 3 revisits the overflow problem in encrypting dynamic controllers. Section 4 introduces secret sharing and Beaver triple. Section 5 proposes a protocol for secure two-party computation of dynamic controllers and analyzes its control performance. It also compares the computational and communication complexities of the proposed protocol with those of previous schemes. Section 6 presents system-theoretic and cryptographic improvements for reducing communication complexity. Section 7 examines the efficiency of the proposed protocol by numerical examples. Finally, Section 8 concludes the paper.
[…]
2 Problem Setting
Consider the discrete-time linear time-invariant plant
$x_{p}(t+1)$ $=A_{p}x_{p}(t)+B_{p}u(t),$ (1)
$y(t)$ $=C_{p}x_{p}(t),$
where $t\in\mathbb{N}_{0}$ is the time index, $x_{p}(t)\in\mathbb{R}^{n_{p}}$ is the state, $u(t)\in\mathbb{R}^{m}$ is the input, $y(t)\in\mathbb{R}^{p}$ is the output, and the initial state is $x_{p}(0)=x_{p,0}$. Suppose the plant is controlled by the discrete-time linear time-invariant controller
$x(t+1)$ $=Ax(t)+By(t),$ (2)
$u(t)$ $=Cx(t)+Dy(t),$
where $x(t)\in\mathbb{R}^{n}$ is the controller state, and its initial value is $x(0)=x_{0}$.
Define the set of $k$-bit fixed-point numbers with an $\ell$-bit fractional part as $\mathbb{Q}_{\langle k,\ell\rangle}\coloneqq\{2^{-\ell}z\mid z\in\mathbb{Z}_{\langle k\rangle}\}$, where $\mathbb{Z}_{\langle k\rangle}\coloneqq\mathbb{Z}_{2^{k}}=\{-2^{k-1},\dots,2^{k-1}-1\}$ represents the set of $k$-bit integers. We assume that the controller parameters and the initial state are represented by $k$-bit fixed-point numbers.
Assumption 1
Let $k$ and $\ell$ be positive integers, such that $k>\ell$. The controller parameters and initial state satisfy $A\in\mathbb{Q}_{\langle k,\ell\rangle}^{n\times n}$, $B\in\mathbb{Q}_{\langle k,\ell\rangle}^{n\times p}$, $C\in\mathbb{Q}_{\langle k,\ell\rangle}^{m\times n}$, $D\in\mathbb{Q}_{\langle k,\ell\rangle}^{m\times p}$, and $x_{0}\in\mathbb{Q}_{\langle k,\ell\rangle}^{n}$.
Note that this assumption is reasonable in practice because fixed-point numbers can approximate real-valued controller parameters and initial state with desired precision $k$ and $\ell$.
We also assume that the closed-loop system consisting of (1) and (2),
$\begin{bmatrix}x_{p}(t+1)\\ x(t+1)\end{bmatrix}=\Phi\begin{bmatrix}x_{p}(t)\\ x(t)\end{bmatrix},\ \Phi\coloneqq\begin{bmatrix}A_{p}+B_{p}DC_{p}&B_{p}C\\ BC_{p}&A\end{bmatrix}$, (3)
is stable.
Assumption 2
The matrix $\Phi$ in (3) is Schur stable.
The goal of this study is to design a secure two-party protocol for running a dynamic controller without state decryption and input re-encryption, where the control performance with the protocol is sufficiently close to that of the original controller (2). In other words, we will propose a secure protocol that outputs $\hat{u}(t)$ satisfying, for any given $\epsilon>0$,
$\norm*{u(t)-\hat{u}(t)}\leq\epsilon$ (4)
for all $t\in\mathbb{N}_{0}$.
We adopt a client-aided computation model [23] for two-party computation. In this setting, the client who initiates queries to two parties provides not only the protocol inputs but also auxiliary inputs, although it does not interact with them during the computation. The parties use these auxiliary inputs to execute the protocol efficiently and return the outputs to the client. We also assume peer-to-peer secure channels as a network model, which can be realized by standard symmetric-key encryption. Furthermore, throughout this paper, we assume the following standard assumptions for two-party computation.
Assumption 3
The parties who perform a protocol are semi-honest and do not collude with each other.
This assumption implies that the parties follow our intended protocol but independently attempt to learn private information of the client from their received messages. It also prohibits external adversaries from corrupting or accessing both parties.
3 Dynamic Controllers Over Integers
Our basic idea for designing a protocol for the secure implementation of the dynamic controller (2) is to split it into two randomized controllers. The parameters and states of these controllers are represented by the shares of a secret sharing scheme. Each randomized controller is installed on a distinct server that computes a share of $u(t)$ using the controller with a share of $y(t)$ and returns it to the plant.
To perform controller computation with secret sharing, the controller parameters, states, inputs, and outputs should be encoded into matrices and vectors over integers. This section revisits the well-known overflow problem [24, 14] caused by such encoding and clarifies the nature of the difficulty of encrypting dynamic controllers using homomorphic encryption.
We begin by encoding the controller parameters, the initial state, and the plant output. Assumption 1 results in the encoded parameters and initial state given as
$\bar{A}$ $\coloneqq 2^{\ell}A\in\mathbb{Z}_{\langle k\rangle}^{n\times n},$ $\ \bar{B}$ $\coloneqq 2^{\ell}B\in\mathbb{Z}_{\langle k\rangle}^{n\times p},$ $\ \bar{C}$ $\coloneqq 2^{\ell}C\in\mathbb{Z}_{\langle k\rangle}^{m\times n},$ (5)
$\bar{D}$ $\coloneqq 2^{\ell}D\in\mathbb{Z}_{\langle k\rangle}^{m\times p},$ $\ \bar{x}_{0}$ $\coloneqq 2^{\ell}x_{0}\in\mathbb{Z}_{\langle k\rangle}^{n}.$
Similarly, the output of the plant (1) can be encoded as
$\bar{y}(t)\coloneqq\left\lceil 2^{\ell}y(t)\right\rfloor\in\mathbb{Z}^{p},$
where the rounding is performed in an element-wise manner.
With the encoded parameters, state, and output, one might consider the following controller over integers [11, 12],
$z(t+1)$ $=\bar{A}z(t)+2^{t\ell}\bar{B}\bar{y}(t),$
$v(t)$ $=\bar{C}z(t)+2^{t\ell}\bar{D}\bar{y}(t),$
where $z(t)\in\mathbb{Z}^{n}$, $v(t)\in\mathbb{Z}^{m}$, and $z(0)=\bar{x}_{0}$. Unfortunately, this controller is problematic due to the increase in the bit length of the controller state. The controller state at time $t\in\mathbb{N}$ is given as
$z(t)\approx 2^{(t+1)\ell}\quantity(A^{t}x_{0}+\sum_{s=0}^{t-1}A^{t-1-s}By(s))=2^{(t+1)\ell}x(t),$
if $\bar{y}(t)\approx 2^{\ell}y(t)$. Hence, the number of bits in the controller state increases by $\ell$ at each time step. This implies that the norm of the encoded state $z(t)$ eventually diverges as time $t$ approaches infinity because $\ell>0$. Therefore, the encoded state causes overflow at a finite time step.
To avoid this problem, one can consider another controller with truncation [14] given as
$\bar{x}(t+1)$ $=\left\lceil\quantity(\bar{A}\bar{x}(t)+\bar{B}\bar{y}(t))/2^{\ell}\right\rfloor,$ (6)
$\bar{u}(t)$ $=\bar{C}\bar{x}(t)+\bar{D}\bar{y}(t),$
where $\bar{x}(t)\in\mathbb{Z}^{n}$ and $\bar{u}(t)\in\mathbb{Z}^{m}$ denote the encoded state and input, respectively, and $\bar{x}(0)=\bar{x}_{0}$. In this controller, the least $\ell$ bits of the state update $\bar{A}\bar{x}(t)+\bar{B}\bar{y}(t)$ are truncated by $\lceil\cdot/2^{\ell}\rfloor$. Then, the bit length of the controller state does not increase because
$\bar{x}(t+1)\approx 2^{\ell}\quantity(Ax(t)+By(t))=2^{\ell}x(t+1),$
for a sufficiently large $\ell$. Additionally, in this case, the control input can be recovered using $u(t)\approx 2^{-2\ell}\bar{u}(t)$.
This approximation suggests that truncation solves the problem of controller state overflow, making (6) suitable for encoding the controller (2). However, implementing the encoded controller using homomorphic encryption remains problematic. The difficulty lies in the fact that homomorphic encryption is supposed to be unable to compute the truncation $\lceil\cdot/2^{\ell}\rfloor$ on encrypted data directly [14]. This is a fundamental problem in the encryption of dynamic controllers using homomorphic encryption. We address this problem by constructing an efficient protocol to compute the truncation using secret sharing, further discussed in Section 5.
4 Secret Sharing and Beaver Triple
This section provides a brief review of secret sharing and its arithmetic [17, 25].
4.1 Secret Sharing
A $t$-out-of-$n$ secret sharing scheme splits a secret into $n$ shares, and the share can be recovered by collecting $t$ shares. In this study, we use a standard 22-out-of-22 secret sharing scheme over $\mathbb{Z}_{q}$ with a prime $q$ defined as follows.
Definition 1 (22-out-of-22 secret sharing)
A 22-out-of-22 secret sharing scheme over $\mathbb{Z}_{q}$ is a tuple of polynomial-time algorithms $\mathsf{Share}$ and $\mathsf{Reconst}$ such that:
  • $\llbracket m\rrbracket\leftarrow\mathsf{Share}(m)$: The share generation algorithm takes a secret $m\in\mathbb{Z}_{q}$ as input, randomly choose $r\in\mathbb{Z}_{q}$, and outputs shares $\llbracket m\rrbracket=(\llbracket m\rrbracket_{1},\llbracket m\rrbracket_{2})=(r,m-r\bmod q)$.
  • $m\leftarrow\mathsf{Reconst}(\llbracket m\rrbracket)$: The reconstruction algorithm takes shares $\llbracket m\rrbracket=(\llbracket m\rrbracket_{1},\llbracket m\rrbracket_{2})$ as input and outputs $m=\llbracket m\rrbracket_{1}+\llbracket m\rrbracket_{2}\bmod q$.
Each share $\llbracket m\rrbracket_{i}$, $i\in\{1,2\}$, is uniformly distributed over $\mathbb{Z}_{q}$ and thus provides no information about $m$ to the party $P_{i}$. Nevertheless, any secret in $\mathbb{Z}_{q}$ can be correctly recovered by collecting the shares, i.e.,
$\mathsf{Reconst}(\mathsf{Share}(m))=m$ (P1)
for all $m\in\mathbb{Z}_{q}$. Furthermore, the secret sharing scheme allows the constant addition, constant multiplication, and addition of shares. For $k,x\in\mathbb{Z}_{q}$, the constant addition and multiplication of $k$ and $\llbracket x\rrbracket$ are defined as $\llbracket x\rrbracket+k\coloneqq(\llbracket x\rrbracket_{1}+k\bmod q,\llbracket x\rrbracket_{2})$ and $k\llbracket x\rrbracket\coloneqq(k\llbracket x\rrbracket_{1}\bmod q,k\llbracket x\rrbracket_{2}\bmod q)$, respectively. The addition of $\llbracket x\rrbracket$ and $\llbracket y\rrbracket$ with $y\in\mathbb{Z}_{q}$ is also defined as $\llbracket x\rrbracket+\llbracket y\rrbracket\coloneqq(\llbracket x\rrbracket_{1}+\llbracket y\rrbracket_{1}\bmod q,\llbracket x\rrbracket_{2}+\llbracket y\rrbracket_{2}\bmod q)$. By definition, it holds that
$\llbracket x\rrbracket+k$ $=\llbracket x+k\bmod q\rrbracket,$ (P2)
$k\llbracket x\rrbracket$ $=\llbracket kx\bmod q\rrbracket,$ (P3)
$\llbracket x\rrbracket+\llbracket y\rrbracket$ $=\llbracket x+y\bmod q\rrbracket.$ (P4)
We also define the subtractions, $\llbracket x\rrbracket-k$ and $\llbracket x\rrbracket-\llbracket y\rrbracket$, similar to the additions. The above computations can be performed without interaction, meaning that $x$ and $y$ are kept secret from both the parties.
4.2 Beaver Triple
Unlike addition and subtraction, the multiplication of shares cannot be naturally defined. That is, $(\llbracket x\rrbracket_{1}\llbracket y\rrbracket_{1}\bmod q,\llbracket x\rrbracket_{2}\llbracket y\rrbracket_{2}\bmod q)\neq\llbracket xy\bmod q\rrbracket$ because $xy\bmod q=(\llbracket x\rrbracket_{1}+\llbracket x\rrbracket_{2})(\llbracket y\rrbracket_{1}+\llbracket y\rrbracket_{2})\bmod q=\llbracket x\rrbracket_{1}\llbracket y\rrbracket_{1}+\llbracket x\rrbracket_{1}\llbracket y\rrbracket_{2}+\llbracket x\rrbracket_{2}\llbracket y\rrbracket_{1}+\llbracket x\rrbracket_{2}\llbracket y\rrbracket_{2}\bmod q$. Although the terms $\llbracket x\rrbracket_{1}\llbracket y\rrbracket_{1}$ and $\llbracket x\rrbracket_{2}\llbracket y\rrbracket_{2}$ can be computed locally by each party, the computations of $\llbracket x\rrbracket_{1}\llbracket y\rrbracket_{2}$ and $\llbracket x\rrbracket_{2}\llbracket y\rrbracket_{1}$ require the exchange of shares, thereby compromising privacy.
To compute valid shares of $xy\bmod q$ privately, we use a technique based on Beaver triples [25]. A Beaver triple over $\mathbb{Z}_{q}$ is a triplet $(\llbracket a\rrbracket,\llbracket b\rrbracket,\llbracket c\rrbracket)$ that consists of shares of random numbers $a,b,c\in\mathbb{Z}_{q}$ such that $c=ab\bmod q$. With the triplet, the multiplication of shares is computed as follows [25, 17]. Suppose that the parties $P_{i}$, $i\in\{1,2\}$ have $\llbracket x\rrbracket_{i}$, $\llbracket y\rrbracket_{i}$, and $(\llbracket a\rrbracket_{i},\llbracket b\rrbracket_{i},\llbracket c\rrbracket_{i})$. Each party computes $\llbracket d\rrbracket_{i}=\llbracket x\rrbracket_{i}-\llbracket a\rrbracket_{i}\bmod q$ and $\llbracket e\rrbracket_{i}=\llbracket y\rrbracket_{i}-\llbracket b\rrbracket_{i}\bmod q$ locally. The parties open $\llbracket d\rrbracket_{i}$ and $\llbracket e\rrbracket_{i}$ and compute $d=\mathsf{Reconst}(\llbracket d\rrbracket)$ and $e=\mathsf{Reconst}(\llbracket e\rrbracket)$, where $\llbracket d\rrbracket=(\llbracket d\rrbracket_{1},\llbracket d\rrbracket_{2})$ and $\llbracket e\rrbracket=(\llbracket e\rrbracket_{1},\llbracket e\rrbracket_{2})$. Then, $P_{1}$ and $P_{2}$ respectively compute $\llbracket z\rrbracket_{1}=e\llbracket a\rrbracket_{1}+d\llbracket b\rrbracket_{1}+\llbracket c\rrbracket_{1}+de\bmod q$ and $\llbracket z\rrbracket_{2}=e\llbracket a\rrbracket_{2}+d\llbracket b\rrbracket_{2}+\llbracket c\rrbracket_{2}\bmod q$ locally. The shares $\llbracket z\rrbracket=(\llbracket z\rrbracket_{1},\llbracket z\rrbracket_{2})$ satisfy $\llbracket z\rrbracket=\llbracket xy\bmod q\rrbracket$ because
$\mathsf{Reconst}(\llbracket z\rrbracket)=\llbracket z\rrbracket_{1}+\llbracket z\rrbracket_{2}\bmod q,$
$=e(\llbracket a\rrbracket_{1}+\llbracket a\rrbracket_{2})+d(\llbracket b\rrbracket_{1}+\llbracket b\rrbracket_{2})+(\llbracket c\rrbracket_{1}+\llbracket c\rrbracket_{2})$
$\phantom{{}={}}+de\bmod q,$
$=(y-b)a+(x-a)b+ab+(x-a)(y-b)\bmod q,$
$=xy\bmod q,$
where we used the properties from (P1) to (P4).
Protocol 1 summarizes the above procedure for computing the multiplication of shares. Using the $\mathsf{Mult}$ protocol, we define
$\llbracket x\rrbracket\llbracket y\rrbracket\coloneqq\mathsf{Mult}(\llbracket x\rrbracket,\llbracket y\rrbracket)=\llbracket xy\bmod q\rrbracket.$ (P5)
Note that the protocol does not leak any information about $x$ and $y$ to the parties because of the randomness of $d=x-a\bmod q$ and $e=y-b\bmod q$.
Protocol 1 Multiplication ($\mathsf{Mult}$)
1:Shares $\llbracket x\rrbracket$ and $\llbracket y\rrbracket$ over $\mathbb{Z}_{q}$ and a Beaver triple $(\llbracket a\rrbracket,\llbracket b\rrbracket,\llbracket c\rrbracket)$ over $\mathbb{Z}_{q}$.
2:Shares $\llbracket z\rrbracket$ over $\mathbb{Z}_{q}$ such that $\llbracket z\rrbracket=\llbracket xy\bmod q\rrbracket$.
[…]
For $M\in\mathbb{Z}_{q}^{d_{1}\times d_{2}}$, denote $\llbracket M\rrbracket\leftarrow\mathsf{Share}(M)$ as a matrix in which the shares of each element of $M$ are arranged in the same manner, namely
$\llbracket M\rrbracket=\begin{bmatrix}\llbracket M_{11}\rrbracket&\cdots&\llbracket M_{1d_{2}}\rrbracket\\ \vdots&\ddots&\vdots\\ \llbracket M_{d_{1}1}\rrbracket&\cdots&\llbracket M_{d_{1}d_{2}}\rrbracket\\ \end{bmatrix}$.
Using the notation and properties (P4) and (P5), the valid shares of matrix addition $X+Y$ for $X,Y\in\mathbb{Z}_{q}^{d_{1}\times d_{2}}$ and multiplication $XZ$ for $X$ and $Z\in\mathbb{Z}_{q}^{d_{2}\times d_{3}}$ can be computed as $\llbracket X\rrbracket+\llbracket Y\rrbracket$ and $\llbracket X\rrbracket\llbracket Z\rrbracket$, respectively. The $(i,j)$ elements of the resultant matrices are $\quantity(\llbracket X\rrbracket+\llbracket Y\rrbracket)_{ij}=\llbracket X_{ij}+Y_{ij}\bmod q\rrbracket$ and $\quantity(\llbracket X\rrbracket\llbracket Z\rrbracket)_{ij}=\llbracket X_{i1}Z_{1j}+\cdots+X_{id_{2}}Z_{d_{2}j}\bmod q\rrbracket$. Here, the matrix multiplication requires distinct $d_{1}d_{2}d_{3}$ Beaver triples. In what follows, the $i$th shares of $\llbracket M\rrbracket$ are denoted as
$\llbracket M\rrbracket_{i}=\begin{bmatrix}\llbracket M_{11}\rrbracket_{i}&\cdots&\llbracket M_{1d_{2}}\rrbracket_{i}\\ \vdots&\ddots&\vdots\\ \llbracket M_{d_{1}1}\rrbracket_{i}&\cdots&\llbracket M_{d_{1}d_{2}}\rrbracket_{i}\\ \end{bmatrix}$
for simplicity of notation. In addition, the reconstruction algorithm $\mathsf{Reconst}$ is supposed to be performed for each element of the matrix, namely $\mathsf{Reconst}(\llbracket M\rrbracket)=M$.
Remark 1
Under Assumption 3, the 22-out-of-22 secret sharing scheme satisfies perfect security [26]. This security means that parties can learn nothing from their own shares, even if their computational capabilities are unlimited.
5 Two-Party Computation of Dynamic Controllers
In this section, we propose an implementation of the encoded controller (6) over the two-party configuration shown in Fig. 11(b) using the 22-out-of-22 secret sharing scheme. The computation of (6) involves addition, multiplication, and truncation. The addition and multiplication of shares are supported by the arithmetic of the secret sharing scheme, (P4) and (P5), introduced in the previous section. However, the truncation of shares cannot be computed directly from (P2) to (P5).
To enable this, we first introduce a truncation protocol originally proposed in [27] and slightly modified here to suit our setting, along with its corresponding security notion. We then apply the protocol to realize (6) over shares and analyze the control performance achievable by the proposed protocol.
5.1 Truncation Protocol
An important insight behind the truncation protocol in [27] is that, by definition of the modulo operation, the truncation of $\ell$ bits of $m\in\mathbb{Z}$ can be represented by
$\left\lceil\frac{m}{2^{\ell}}\right\rfloor$ $=\frac{m-(m\bmod 2^{\ell})}{2^{\ell}},$
$=\operatorname{inv}(2^{\ell},q)(m-(m\bmod 2^{\ell}))\bmod q,$
if $q$ is a prime such that $|q/2>\absolutevalue*{\lceil m/2^{\ell}\rfloor}$ and $q/2>2^{\ell}$. Notice that $\operatorname{inv}(2^{\ell},q)$ is the modular inverse of $2^{\ell}$ modulo $q$ and can be computed using the extended Euclidean algorithm. Here, let $q$ be the prime modulus of the secret sharing scheme in the previous section, and $\lambda\in\mathbb{N}$ be an appropriately chosen security parameter. Suppose that the bit length of a message $m$ is at most $\kappa=\lfloor\log_{2}q\rfloor-\lambda-1$ and $\ell$ is less than $\kappa$. To simplify our protocol, we assume that $\ell$ is public, which is acceptable because it specifies only the bit length of fractional part. Based on the above observation, the $\mathsf{Trunc}$ protocol shown in Protocol 2 computes the truncation of $\ell$ bits of a message $m\in\mathbb{Z}_{\langle\kappa\rangle}$ over shares.
Protocol 2 Truncation ($\mathsf{Trunc}$)
1:Shares $\llbracket m\rrbracket$ over $\mathbb{Z}_{q}$ with $m\in\mathbb{Z}_{\langle\kappa\rangle}$, shares $(\llbracket r\rrbracket,\llbracket r^{\prime}\rrbracket)$ over $\mathbb{Z}_{q}$ with random numbers $r\in\mathbb{Z}_{\langle\kappa-\ell+\lambda\rangle}$ and $r^{\prime}\in\mathbb{Z}_{\langle\ell\rangle}$, and bit length $\ell\in\mathbb{N}$, where $\kappa=\lfloor\log_{2}q\rfloor-\lambda-1>\ell$, and $\lambda\in\mathbb{N}$ is a security parameter.
2:Shares $\llbracket m^{\prime}\rrbracket$ over $\mathbb{Z}_{q}$ such that $m^{\prime}=\lceil m/2^{\ell}\rfloor+w$, where $w\in\{-1,0,1\}$.
[…]
The protocol takes as input not only $\llbracket m\rrbracket=(\llbracket m\rrbracket_{1},\llbracket m\rrbracket_{2})$ and $\ell$ but also shares $\llbracket r\rrbracket=(\llbracket r\rrbracket_{1},\llbracket r\rrbracket_{2})$ and $\llbracket r^{\prime}\rrbracket=(\llbracket r^{\prime}\rrbracket_{1},\llbracket r^{\prime}\rrbracket_{2})$ with random numbers $r\in\mathbb{Z}_{\langle\kappa-\ell+\lambda\rangle}$ and $r^{\prime}\in\mathbb{Z}_{\langle\ell\rangle}$. Suppose that parties $P_{i}$, $i\in\{1,2\}$ have $\llbracket m\rrbracket_{i}$, $\llbracket r\rrbracket_{i}$, and $\llbracket r^{\prime}\rrbracket_{i}$. Each party computes $\llbracket m_{r}\rrbracket_{i}=\llbracket m\rrbracket_{i}+2^{\ell}\llbracket r\rrbracket_{i}+\llbracket r^{\prime}\rrbracket_{i}+2^{\ell-1}\bmod q$ locally. The party $P_{2}$ transmits $\llbracket m_{r}\rrbracket_{2}$ to $P_{1}$, and $P_{1}$ reconstructs $m_{r}=\mathsf{Reconst}(\llbracket m_{r}\rrbracket)$, where $\llbracket m_{r}\rrbracket=(\llbracket m_{r}\rrbracket_{1},\llbracket m_{r}\rrbracket_{2})$. Then, $P_{1}$ and $P_{2}$ locally compute $\llbracket m^{\prime}\rrbracket_{1}=\operatorname{inv}(2^{\ell},q)(\llbracket m\rrbracket_{1}+\llbracket r^{\prime}\rrbracket_{1}-(m_{r}-2^{\ell-1}\bmod 2^{\ell}))\bmod q$ and $\llbracket m^{\prime}\rrbracket_{2}=\operatorname{inv}(2^{\ell},q)(\llbracket m\rrbracket_{2}+\llbracket r^{\prime}\rrbracket_{2})\bmod q$, respectively. Consequently, $\llbracket m^{\prime}\rrbracket=(\llbracket m^{\prime}\rrbracket_{1},\llbracket m^{\prime}\rrbracket_{2})$ becomes shares of $\lceil m/2^{\ell}\rfloor+w$ for all $m\in\mathbb{Z}_{\langle\kappa\rangle}$, where $w\in\{-1,0,1\}$.
Lemma 1
Suppose $q$ is a prime. Let $\kappa,\ell,\lambda\in\mathbb{N}$ such that $\kappa=\lfloor\log_{2}q\rfloor-\lambda-1>\ell$. It holds that, for all $m\in\mathbb{Z}_{\langle\kappa\rangle}$,
$\mathsf{Reconst}\quantity(\mathsf{Trunc}\quantity(\llbracket m\rrbracket,\ell))=\left\lceil\frac{m}{2^{\ell}}\right\rfloor+w,$
where $\llbracket m\rrbracket\leftarrow\mathsf{Share}(m)$ and $w\in\{-1,0,1\}$.
Proof 5.1.
See Appendix .1.
The lemma shows that the $\mathsf{Trunc}$ protocol truncates the least $\ell$ bits of $m\in\mathbb{Z}_{\langle\kappa\rangle}$ with a $1$-bit error $w$. This error occurs only when $m+r^{\prime}\notin\mathbb{Z}_{\langle\ell\rangle}$ and thus can be removed by an additional protocol to detect $m+r^{\prime}\notin\mathbb{Z}_{\langle\ell\rangle}$ over shares [27]. The following section tolerates the error to avoid increasing the computational and communication costs and applies the truncation protocol to approximate and implement the encoded controller (6) to run it over shares.
Remark 5.2.
The security of $\mathsf{Trunc}$ is identical to that of the truncation protocol in [27]. Through protocol execution, $P_{2}$ does not receive any messages from $P_{1}$ and thus learns no information. The only information leaked into $P_{1}$ is $m_{r}=m+2^{\ell}r+r^{\prime}+2^{\ell-1}\bmod q$. Under the conditions of bit lengths of $m$, $r$, and $r^{\prime}$, it holds that $m_{r}=m+r^{\prime\prime}\in\mathbb{Z}_{q}$, where $r^{\prime\prime}=2^{\ell}r+r^{\prime}+2^{\ell-1}$ is uniformly distributed over $\mathbb{Z}_{\langle\kappa+\lambda\rangle}$. Intuitively, $m_{r}$ is a masked message of $m$ by $r^{\prime\prime}$, and thus $P_{1}$ cannot learn any information from $m_{r}$. Indeed, the statistical distance between $m+r^{\prime\prime}$ and $r^{\prime\prime}$, $|\frac{1}{2}\sum_{x\in\mathbb{Z}_{q}}\absolutevalue*{\Pr[m+r^{\prime\prime}=x]-\Pr[r^{\prime\prime}=x]}$, is less than $2^{-\lambda}$ and so is negligible in $\lambda$. Therefore, the $\mathsf{Trunc}$ protocol achieves statistical security, which is a relaxation of the perfect security. However, statistical security is still valid for unconditional adversaries. This implies that it is stronger than computational security achieved by homomorphic encryption [5, 6, 7, 8, 9]. Note that computational security covers only conditional adversaries whose computational capability is limited to polynomial time [28].
5.2 Controller Computation Protocol
With Protocol 1 and Protocol 2, we propose Protocol 3, which enables running a dynamic controller over the secret sharing scheme without state decryption and input re-encryption. For convenience, we define the plant controlled by the proposed protocol instead of the controller (2) as
$\hat{x}_{p}(t+1)$ $=A_{p}\hat{x}_{p}(t)+B_{p}\hat{u}(t),$ (7)
$\hat{y}(t)$ $=C_{p}\hat{x}_{p}(t),$
where $\hat{x}_{p}(t)\in\mathbb{R}^{n_{p}}$ is the state, $\hat{u}(t)\in\mathbb{R}^{m}$ is the input provided by the protocol, $\hat{y}(t)\in\mathbb{R}^{p}$ is the output, and $\hat{x}_{p}(0)=x_{p}(0)=x_{p,0}$ is the initial state. In that case, the encoded plant output is redefined as
$\bar{y}(t)\coloneqq\left\lceil 2^{\ell}\hat{y}(t)\right\rfloor\in\mathbb{Z}^{p}.$ (8)
In what follows, we refer to the plant (7) with a sensor and actuator as a client and two servers computing a dynamic controller as parties $P_{1}$ and $P_{2}$.
Protocol 3 Client-aided two-party controller computation
1:Controller parameters $A\in\mathbb{Q}_{\langle k,\ell\rangle}^{n\times n}$, $B\in\mathbb{Q}_{\langle k,\ell\rangle}^{n\times p}$, $C\in\mathbb{Q}_{\langle k,\ell\rangle}^{m\times n}$, $D\in\mathbb{Q}_{\langle k,\ell\rangle}^{m\times p}$, initial state $x_{0}\in\mathbb{Q}_{\langle k,\ell\rangle}^{n}$, plant output $\hat{y}(t)\in\mathbb{R}^{n}$, bit lengths $k,\ell\in\mathbb{N}$, and modulus $q$
2:Control input $\hat{u}(t)\in\mathbb{R}^{m}$
[…]
4:C: Compute $\big{\llbracket}\tilde{A}\big{\rrbracket}$, $\big{\llbracket}\tilde{B}\big{\rrbracket}$, $\big{\llbracket}\tilde{C}\big{\rrbracket}$, $\big{\llbracket}\tilde{D}\big{\rrbracket}$, and $\big{\llbracket}\tilde{x}(0)\big{\rrbracket}$ by (9)
5:C: Send $\big{\llbracket}\tilde{A}\big{\rrbracket}_{i}$, $\big{\llbracket}\tilde{B}\big{\rrbracket}_{i}$, $\big{\llbracket}\tilde{C}\big{\rrbracket}_{i}$, $\big{\llbracket}\tilde{D}\big{\rrbracket}_{i}$, and $\big{\llbracket}\tilde{x}(0)\big{\rrbracket}_{i}$ to $P_{i}$
[…]
7:for all $t\in\mathbb{N}_{0}$
8: C: Compute $\llbracket\tilde{y}(t)\rrbracket$ by (10)
9: C: Generate auxiliary inputs $\mathsf{aux}_{1}$ and $\mathsf{aux}_{2}$
10: C: Send $\llbracket\tilde{y}(t)\rrbracket_{i}$ and $\mathsf{aux}_{i}$ to $P_{i}$
[…]
13: C: Compute $\hat{u}(t)$ by (13)
To implement (6) with the secret sharing scheme, the client splits its controller parameters and initial state into shares using the share generation algorithm,
[…]
where
$\tilde{A}$ $\coloneqq\bar{A}\bmod q,$ $\quad\tilde{B}$ $\coloneqq\bar{B}\bmod q,$ $\quad\tilde{C}$ $\coloneqq\bar{C}\bmod q,$
$\tilde{D}$ $\coloneqq\bar{D}\bmod q,$ $\quad\tilde{x}_{0}$ $\coloneqq\bar{x}_{0}\bmod q,$
and $\bar{A}$, $\bar{B}$, $\bar{C}$, $\bar{D}$, and $\bar{x}_{0}$ are the encoded controller parameters and initial state defined in (5). In the offline phase, as shown in Fig. 22(a), the client sends the $i$th shares of the parameters $\big{\llbracket}\tilde{A}\big{\rrbracket}_{i},\big{\llbracket}\tilde{B}\big{\rrbracket}_{i},\big{\llbracket}\tilde{C}\big{\rrbracket}_{i},\big{\llbracket}\tilde{D}\big{\rrbracket}_{i}$ and initial state $\big{\llbracket}\tilde{x}(0)\big{\rrbracket}_{i}$ to the party $P_{i}$. The parties store the received shares individually and use them in the online phase later.
In the online phase, the parties run an approximate dynamic controller for (6) as shown in Fig. 22(b). For each time $t\in\mathbb{N}_{0}$, the client reads the plant output $\hat{y}(t)$ in (7) and generates its shares by
$\llbracket\tilde{y}(t)\rrbracket\leftarrow\mathsf{Share}\quantity(\tilde{y}(t)),\quad\tilde{y}(t)=\bar{y}(t)\bmod q\in\mathbb{Z}_{q}^{p},$ (10)
where $\bar{y}(t)$ is the encoded output defined in (8). The $i$th shares of the output $\llbracket\tilde{y}(t)\rrbracket_{i}$ are sent to the party $P_{i}$. Simultaneously, the client generates $(n+m)(n+p)$ Beaver triples $(\llbracket a_{j}\rrbracket,\llbracket b_{j}\rrbracket,\llbracket c_{j}\rrbracket)$ over $\mathbb{Z}_{q}$ such that $c_{j}=a_{j}b_{j}\bmod q$ for every $j=1,\dots,(n+m)(n+p)$ and shares $(\llbracket r_{h}\rrbracket,\llbracket r^{\prime}_{h}\rrbracket)$ over $\mathbb{Z}_{q}$ with random numbers $r_{h}\in\mathbb{Z}_{\langle\kappa-\ell+\lambda\rangle}$ and $r^{\prime}_{h}\in\mathbb{Z}_{\langle\ell\rangle}$ for every $h=1,\dots,n$. The client then sends the $i$th shares
$\mathsf{aux}_{i}$ $\coloneqq\{(\llbracket a_{j}\rrbracket_{i},\llbracket b_{j}\rrbracket_{i},\llbracket c_{j}\rrbracket_{i}),(\llbracket r_{h}\rrbracket_{i},\llbracket r^{\prime}_{h}\rrbracket_{i})$ (11)
$\phantom{{}\coloneqq{}}\mid j=1,\dots,(n+m)(n+p),h=1,\dots,n\}$
to $P_{i}$ as auxiliary inputs. The parties compute the shares of state update $\tilde{x}(t+1)\in\mathbb{Z}^{n}$ and encoded input $\tilde{u}(t)\in\mathbb{Z}^{m}$ as
$\big{\llbracket}\tilde{x}(t+1)\big{\rrbracket}$ $\leftarrow\mathsf{Trunc}\quantity(\big{\llbracket}\tilde{A}\big{\rrbracket}\big{\llbracket}\tilde{x}(t)\big{\rrbracket}+\big{\llbracket}\tilde{B}\big{\rrbracket}\big{\llbracket}\tilde{y}(t)\big{\rrbracket},\ell),$ (12)
$\big{\llbracket}\tilde{u}(t)\big{\rrbracket}$ $=\big{\llbracket}\tilde{C}\big{\rrbracket}\big{\llbracket}\tilde{x}(t)\big{\rrbracket}+\big{\llbracket}\tilde{D}\big{\rrbracket}\big{\llbracket}\tilde{y}(t)\big{\rrbracket},$
where the $\mathsf{Trunc}$ protocol is applied for each element of $\big{\llbracket}\tilde{A}\big{\rrbracket}\big{\llbracket}\tilde{x}(t)\big{\rrbracket}+\big{\llbracket}\tilde{B}\big{\rrbracket}\big{\llbracket}\tilde{y}(t)\big{\rrbracket}$. Note that the multiplications of shares for the state and input computations consume $n(n+p)$ and $m(n+p)$ Beaver triples, respectively. Each party $P_{i}$ returns its shares of the control input $\llbracket\tilde{u}(t)\rrbracket_{i}$ to the client separately. Upon receiving the shares, the client reconstructs and decodes an approximate control input by
$\hat{u}(t)=2^{-2\ell}\tilde{u}(t),\quad\tilde{u}(t)=\mathsf{Reconst}\quantity(\llbracket\tilde{u}(t)\rrbracket).$ (13)
(a) Offline phase.
(b) Online phase.
Figure 2: Client-aided two-party dynamic controller computation.
We now proceed to show the main results of this study. The lemma below guarantees that the proposed controller (12) can operate for an infinite time horizon without causing state overflow and does not require state decryption or input re-encryption.
Lemma 5.3.
Consider the plant (7) and controller (12) under Assumption 1 and Assumption 2. Define $\Gamma\coloneqq[(B_{p}D)^{\top}\ B^{\top}]^{\top}$, $\alpha\coloneqq\norm{C_{p}}_{\infty}+3/2$, and $\beta\coloneqq 2^{\ell}\norm{\chi_{0}}_{\infty}+\norm{\Gamma}_{\infty}/2+3/2$, where $\chi_{0}=[x_{p,0}^{\top}\ x_{0}^{\top}]^{\top}$. If $q$ is a prime, and
$\log_{2}q>k+\lambda+2+\left\lfloor\log_{2}\max\{n,p\}\alpha\beta\frac{c}{1-\gamma}\right\rfloor,$ (14)
then it holds that
$\tilde{x}(t+1)$ $=\left\lceil\quantity(\bar{A}\tilde{x}(t)+\bar{B}\bar{y}(t))/2^{\ell}\right\rfloor+w(t),\quad\tilde{x}(0)=\bar{x}_{0},$
$\tilde{u}(t)$ $=\bar{C}\tilde{x}(t)+\bar{D}\bar{y}(t),$
for all $t\in\mathbb{N}_{0}$, where $\tilde{x}(t)=\mathsf{Reconst}\quantity(\llbracket\tilde{x}(t)\rrbracket)$, $\tilde{u}(t)=\mathsf{Reconst}\quantity(\llbracket\tilde{u}(t)\rrbracket)$, $w(t)\in\{-1,0,1\}^{n}$, and $c\geq 1$ and $\gamma\in(0,1)$ are constants such that $\norm*{\Phi^{t}}\leq c\gamma^{t}$ for all $t\in\mathbb{N}_{0}$.
Proof 5.4.
See Appendix .2.
Lemma 5.3 implies that the reconstruction of the proposed controller (12) is equivalent to the encoded controller (6) with the error $w(t)$ induced by the $\mathsf{Trunc}$ protocol. Based on this result, the following theorem states that the control performance of the proposed controller can be arbitrarily close to that of the original controller (2) by choosing the bit length of the fractional part $\ell$ to be sufficiently large.
Theorem 5.5.
Consider the plants (1) and (7) and controllers (2) and (12) under Assumption 1 and Assumption 2. Suppose (14) is satisfied. For every $\epsilon>0$, if
$\ell\geq\log_{2}\frac{c\epsilon^{-1}}{1-\gamma}\quantity(\frac{\sqrt{p}}{2}\quantity(\norm{\Gamma}\norm{\Upsilon}+\norm{D})+2\sqrt{n}\norm{\Upsilon}),$ (15)
the controllers guarantee (4) for all $t\in\mathbb{N}_{0}$, where $\Upsilon\coloneqq[DC_{p}\ C]$, and $\Gamma$, $c$, and $\gamma$ are defined in Lemma 5.3.
Proof 5.6.
See Appendix .3.
Furthermore, if there is no direct feedthrough in the controllers, a more concise condition can be obtained.
Corollary 5.7.
Suppose $D=\mathbf{0}_{m\times p}$. If
$\ell\geq k-\ell+\log_{2}\frac{c\epsilon^{-1}}{1-\gamma}\quantity(mp\sqrt{n}+n\sqrt{m}),$ (16)
the condition (15) is satisfied.
Proof 5.8.
See Appendix .4.
Here, $k$-$\ell$ in (16) represents the bit length of the integer part of the controller parameters and initial state. This result suggests the following procedure for selecting the parameters used in the proposed controller computation protocol: 1) Given the design parameter $\epsilon>0$ and the security parameter $\lambda\in\mathbb{N}$. 2) Fix $k$-$\ell\in\mathbb{N}$ to be sufficiently large. 3) Choose the bit length $\ell\in\mathbb{N}$. 4) Design the controller (2) to satisfy Assumption 1 and Assumption 2. 5) When $D\neq\mathbf{0}_{m\times p}$, check the condition (15). Otherwise, check the condition (16). If $\ell$ does not satisfy the condition, increase $\ell$ and return to step 4). 6) Choose the modulus $q$ to satisfy (14).
Note that the term $k+\lambda+2$ in (14) is essential to guarantee the correctness and security of the $\mathsf{Trunc}$ protocol. The last term of (14) is induced by encoding and mainly depends on $2^{\ell}$ in $\beta$. This implies a trade-off between communication costs and precision. This trade-off can be improved by optimizing the lower bound of $\ell$ in (15).
Remark 5.9.
During the execution of Protocol 3, each party $P_{i}$ receives the $i$th shares in (9), (10), and (11) from the client and can access public parameters $k$, $\ell$, $q$, and $\lambda$. It also obtains the masked messages $(d,e)$ and $m_{r}$ (if $i=1$) for each invocation of $\mathsf{Mult}$ and $\mathsf{Trunc}$, respectively. Here, the subprotocols are sequentially invoked in the online phase. In this case, the modular composition theorem in [29] guarantees that the security of the entire protocol is equivalent to the weakest security of the subprotocols. Therefore, the proposed protocol achieves statistical security with respect to $\lambda$.
5.3 Comparison to Previous Works
We compare the theoretical costs in the online phase of the proposed protocol with conventional single-server schemes based on learning with errors (LWE) encryption [14] and ring learning with errors (RLWE) encryption [15, 16]. Table 1 lists their computational complexities on the client and server sides, as well as communication complexities from the client (resp. server) to the server (resp. client) and between the servers, which are detailed below. For a fair comparison, the moduli of the ciphertext spaces and the dimensions of the secret keys in the encryption schemes were set to $q$ and $N$, respectively.
Table 1: Comparison of computational and communication complexities
Computational complexity Communication complexity
Scheme Client Server Client $\rightarrow$ Server Server $\rightarrow$ Client Server $\leftrightarrow$ Server
[14] $O((m+p)N)$ $O((n+m)(n+p)\zeta N^{2})$ $O((m+p)N\log q)$ $O(mN\log q)$
[15] $O(N\log N)$ $O((n+m+p)\zeta N\log N)$ $O(N\log q)$ $O(N\log q)$
[16] $O(m^{2}+N\log N)$ $O(nN\log N)$ $O(N\log q)$ $O(N\log q)$
Protocol 3 $O((n+m)(n+p))$ $O((n+m)(n+p))$ $O((n+m)(n+p)\log q)$ $O(m\log q)$ $O((n+m)(n+p)\log q)$
5.3.1 LWE-based encrypted controller in [14]
The authors of [14] represented the controller (2) in the form
$x^{\prime}(t+1)$ $=T(A-RC)T^{-1}x^{\prime}(t)+T(B-RD)y(t)+\mu(t),$
$u(t)$ $=CT^{-1}x^{\prime}(t)+Dy(t),$
with $R\in\mathbb{R}^{n\times m}$, $T\in\mathbb{R}^{n\times n}$, $x^{\prime}(t)=Tx(t)\in\mathbb{R}^{n}$, and $\mu(t)=TRu(t)\in\mathbb{R}^{m}$ such that $T(A-RC)T^{-1}\in\mathbb{Z}^{n\times n}$. They encrypted each element of its parameters and signals using the Gentry-Sahai-Waters (GSW) [8] and LWE [5] encryption schemes, respectively. The ciphertexts of the GSW and LWE encryptions are a matrix in $\mathbb{Z}_{q}^{(N+1)\times\zeta(N+1)}$ and a vector in $\mathbb{Z}_{q}^{N+1}$, where $\zeta=\lfloor\log_{\nu}q\rfloor$ for some $\nu\ll q$.
Communication complexity: For each time $t\in\mathbb{N}_{0}$, the client sends and receives the LWE ciphertexts of $y(t)$ and $u(t)$, respectively, and then returns the re-encryption of $u(t)$ for computing $\mu(t)$. Thus, the communication complexities from the client to the server and from the server to the client are $O((m+p)N\log q)$ and $O(mN\log q)$, respectively.
Computational complexity: The computational complexity of the client is $O((m+p)N)$ because LWE encryption and decryption require an inner product of $N$-dimensional vectors.
On the server, the encrypted controller requires $n(n+2m+p)+mp$ multiplications between the GSW and LWE ciphertexts and n(n+2m+p4)+m(p1)n(n+2m+p-4)+m(p-1) additions of the LWE ciphertexts. According to [16], each GSW-LWE multiplication comprises $\zeta(N+1)^{2}$ scalar multiplications. In addition, each LWE addition is an $(N+1)$-dimensional vector addition. Consequently, the computational complexity of the server is $O((n+m)(n+p)\zeta N^{2})$.
5.3.2 RLWE-based encrypted controller in [15]
The authors of [15] used the controller representation $u(t)=Hd(t)$ with $d(t)=[y(t-n)^{\top}\ \cdots\ y(t)^{\top}\ u(t-n)^{\top}\ \cdots\ u(t-1)^{\top}]^{\top}$ and encrypted $H\in\mathbb{R}^{m\times(n(m+p)+p)}$ and $d(t)\in\mathbb{R}^{n(m+p)+p}$ using the RLWE encryption [7]. The ciphertext of the RLWE encryption is a pair of polynomials in $R_{q}^{2}$, where $R_{q}\coloneqq\mathbb{Z}_{q}[X]/(X^{N}+1)$ is a polynomial ring consisting of $N$-dimensional polynomials in $X$ over $\mathbb{Z}_{q}$.
Communication complexity: Similar to the LWE-based scheme, for each time $t\in\mathbb{N}_{0}$, the client sends the encryption of $y(t)$ and re-encryption of $u(t)$ and receives the encryption of $u(t)$. Meanwhile, the RLWE-based scheme applied message packing [30] when encrypting $y(t)$ and $u(t)$, which enables the encryption of a vector as a single ciphertext in $R_{q}^{2}$. Therefore, the communication complexity from the client (resp. server) to the server (resp. client) is $O(N\log q)$.
Computational complexity: The message packing and unpacking are respectively performed before encryption and after decryption. The packing and unpacking are implemented by the (inverse) number theoretic transform [15], of which complexity is $O(N\log N)$ [31]. The encryption and decryption include a multiplication of polynomials in $R_{q}$. The $N$-dimensional polynomial multiplication can also be computed using the number theoretic transform. Therefore, the computational complexity of the client is $O(N\log N)$.
The encrypted controller requires $n+1$ multiplications with linearization, $n+m+p+2$ additions, and m+p1m+p-1 rotations of the RLWE ciphertexts. Note that the multiplication in the RLWE encryption increases the ciphertext size from $R_{q}^{2}$ to $R_{q}^{3}$. The linearization reduces it to $R_{q}^{2}$, and the rotation replaces the position of vector elements packed in a ciphertext. Each multiplication and addition process includes four multiplications and two additions in $N$-dimensional polynomials, respectively. The linearization and rotation carry out $2\zeta$ NN-dimensional polynomial multiplications [16], where $\zeta$ is as in the LWE-based scheme. Therefore, the computational complexity of the server is $O((n+m+p)\zeta N\log N)$.
5.3.3 RLWE-based encrypted controller in [16]
When $D=\mathbf{0}_{m\times p}$, the authors of [16] employed a variant of the controller representation in [15], $u(t)=H^{\prime}d^{\prime}(t)$, $d^{\prime}(t)=[y(t-1)^{\top}\ \cdots\ y(t-n)^{\top}\ u(t-1)^{\top}\ \cdots\ u(t-n)^{\top}]^{\top}$, and encrypted $H^{\prime}\in\mathbb{R}^{m\times n(m+p)}$ and $d^{\prime}(t)\in\mathbb{R}^{n(m+p)}$ using the RLWE encryption [6].
Communication complexity: The communication complexities are the same as those in [15].
Computational complexity: This RLWE-based scheme considers a different encoding of data and replaces the addition, linearization, and rotation on the server in [15] with aggregation on the client, which consists of $m^{2}$ scalar additions. This changes the computational complexities of the client and server to $O(m^{2}+N\log N)$ and $O(nN\log N)$, respectively.
5.3.4 Proposed protocol (Protocol 3)
First, we note that the computational complexities of the $\mathsf{Share}$ and $\mathsf{Reconst}$ algorithms are $O(1)$ for a scalar message. Thus, the complexities of (constant) addition/multiplication and truncation of shares are $O(1)$.
Communication complexity: For each time $t\in\mathbb{N}_{0}$, the client transmits $\llbracket\tilde{y}(t)\rrbracket_{i}$ in (10) and $\mathsf{aux}_{i}$ in (11) to each party $P_{i}$, $i\in\{1,2\}$ and receives $\llbracket\tilde{u}(t)\rrbracket_{i}$ in (12) from the party. Hence, the communication complexities from the client to the servers and from the servers to the client are $O((n+m)(n+p)\log q)$ and $O(m\log q)$, respectively. Furthermore, to compute the controller (12), the parties invoke the $\mathsf{Mult}$ and $\mathsf{Trunc}$ protocols $(n+m)(n+p)$ and $n$ times, respectively. Each invocation of $\mathsf{Mult}$ and $\mathsf{Trunc}$ requires the transmission of four and one integers in $\mathbb{Z}_{q}$ to exchange $(\llbracket d\rrbracket,\llbracket e\rrbracket)$ and $\llbracket m_{r}\rrbracket_{2}$. Therefore, the communication complexity between the servers is $O((n+m)(n+p)\log q)$.
Computational complexity: The client generates $\llbracket\tilde{y}(t)\rrbracket_{i}$ and $\mathsf{aux}_{i}$ and reconstructs $\llbracket\tilde{u}(t)\rrbracket_{i}$ in the online phase. This results in $(n+m)(n+p)+2n+p$ and $m$ invocations for the $\mathsf{Share}$ and $\mathsf{Reconst}$ algorithms, respectively. In addition to the $\mathsf{Mult}$ and $\mathsf{Trunc}$ protocols, the computation of (12) requires $n+m$ additions of shares. Consequently, the computational complexities of the client and servers are $O((n+m)(n+p))$.
Compared to the previous schemes, the proposed protocol significantly improves the computational efforts of both the client and servers. Note that the dimension parameter $N$ is typically chosen to be larger than $2^{10}$ [32]. On the other hand, although the proposed protocol reduces the communication complexity from the servers to the client, the reduction in the complexity from the client to the servers is limited to satisfying $(n+m)(n+p)\ll N$. This implies that the proposed protocol incurs high communication costs for large-scale controllers.
Remark 5.10.
One might notice that the computational complexity of the client in Protocol 3 is the same as that of the unencrypted case. This is because the client needs to generate the auxiliary inputs $\mathsf{aux}_{i}$, of which size is $O((n+m)(n+p))$. However, the auxiliary inputs are independent of the protocol inputs (e.g., the plant output $\hat{y}(t)$). This implies that their generation can be addressed in the offline phase if the control system terminates in a finite time. Even if this is not the case, it can be outsourced to a third party called a trusted dealer, who is non-colluding with the parties. Then, the computational complexity of the client can be reduced to $O(m+p)$.
Remark 5.11.
Protocol 3 can be extended to $M$ parties using an $M$-out-of-$M$ secret sharing scheme. Then, the complexities in Table 1 increase linearly with $M$. Meanwhile, Assumption 3 can be relaxed to allow the collusion of any M1M-1 parties, which mitigates security risks. Recall that the proposed scheme is compromised when all parties are corrupted, unlike the single-server setting using homomorphic encryption.
6 Improvement of Communication Complexities
As reviewed, the client and parties $P_{i}$, $i\in\{1,2\}$ in the proposed protocol exchange $\llbracket\tilde{y}(t)\rrbracket_{i}$ in (10), $\mathsf{aux}_{i}$ in (11), and $\llbracket\tilde{u}(t)\rrbracket_{i}$ in (12) for each time $t\in\mathbb{N}_{0}$. Thus, the total communication cost between the client and parties is $2(3(n+m)(n+p)+2n+m+p)\ell_{q}$ bit, where $\ell_{q}=\lfloor\log_{2}q\rfloor+1$ is the bit length of the modulus $q$. This section discusses modifications to the proposed protocol, which improves the cost by applying coordinate transformation and cryptographic primitives.
6.1 Brunovsky normal form
It is expected that sparse controller parameters will reduce the number of multiplications of shares by ignoring their zero elements. To transform the parameters of (2) to sparse parameters, we employ the Brunovsky normal form [33].
Definition 6.12 (Brunovsky normal form [34]).
Consider $F\in\mathbb{R}^{n\times n}$ and $G\in\mathbb{R}^{n\times m}$. For $r\leq m$, let $(n_{1},\dots,n_{r})$ be a sequence of positive integers, such that $n_{1}\geq\cdots\geq n_{r}$ and $n_{1}+\cdots+n_{r}=n$. The pair $(F,G)$ is in the Brunovsky normal form if $F$ and $G$ are of the form
$F=\begin{bmatrix}F_{n_{1}}&&\\ &\ddots&\\ &&F_{n_{r}}\end{bmatrix},\quad G=\begin{bmatrix}G_{n_{1}}&&&0&\cdots&0\\ &\ddots&&\vdots&\ddots&\vdots\\ &&G_{n_{r}}&0&\cdots&0\end{bmatrix}$,
where $F_{n_{i}}\in\mathbb{R}^{n_{i}\times n_{i}}$ and $G_{n_{i}}\in\mathbb{R}^{n_{i}\times 1}$ are of the form
$F_{n_{i}}=\begin{bmatrix}0&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\\ f_{n_{i},0}&f_{n_{i},1}&\cdots&f_{n_{i},n_{i}-1}\end{bmatrix},\quad G_{n_{i}}=\begin{bmatrix}0\\ \vdots\\ 0\\ 1\end{bmatrix}$,
for every $i=1,\dots,r$, and $\sum_{j=0}^{n_{i}}f_{n_{i},j}X^{j}$ with $f_{n_{i},n_{i}}=1$ is the characteristic polynomial of $F_{n_{i}}$.
[…]
$\begin{bmatrix}x_{n_{1}}(t+1)\\ \vdots\\ x_{n_{m}}(t+1)\end{bmatrix}$ $\!=\!\underbrace{\begin{bmatrix}A_{n_{1}}&&\\ &\ddots&\\ &&A_{n_{m}}\end{bmatrix}}_{(P^{\top})^{-1}AP^{\top}}\begin{bmatrix}x_{n_{1}}(t)\\ \vdots\\ x_{n_{m}}(t)\end{bmatrix}+\underbrace{\begin{bmatrix}B_{n_{1}}\\ \vdots\\ B_{n_{m}}\end{bmatrix}}_{(P^{\top})^{-1}B}y(t),$
$\begin{bmatrix}u_{1}(t)\\ \vdots\\ u_{m}(t)\end{bmatrix}$ $\!=\!\underbrace{\begin{bmatrix}C_{n_{1}}&&\\ &\ddots&\\ &&C_{n_{m}}\end{bmatrix}}_{CP^{\top}}\begin{bmatrix}x_{n_{1}}(t)\\ \vdots\\ x_{n_{m}}(t)\end{bmatrix}+\underbrace{\begin{bmatrix}D_{1}\\ \vdots\\ D_{m}\end{bmatrix}}_{D}y(t),$
where $A_{n_{i}}\in\mathbb{R}^{n_{i}\times n_{i}}$ and $C_{n_{i}}\in\mathbb{R}^{1\times n_{i}}$ have the same form as $F_{n_{i}}^{\top}$ and $G_{n_{i}}^{\top}$ in Definition 6.12, respectively, $B_{n_{i}}\in\mathbb{R}^{n_{i}\times p}$, and $D_{i}\in\mathbb{R}^{1\times p}$.
We replace $(A,B,C,D)$ of the controller (2) with $((P^{\top})^{-1}AP^{\top},(P^{\top})^{-1}B,CP^{\top},D)$ and assume that the parameters satisfy Assumption 1. Using the decomposition, the proposed controller (12) can be implemented as, for every $i=1,\dots,m$,
$\big{\llbracket}\tilde{x}_{n_{i}}(t+1)\big{\rrbracket}$ $\leftarrow\mathsf{Trunc}\quantity(\big{\llbracket}\tilde{A}_{n_{i}}\big{\rrbracket}\big{\llbracket}\tilde{x}_{n_{i}}(t)\big{\rrbracket}+\big{\llbracket}\tilde{B}_{n_{i}}\big{\rrbracket}\big{\llbracket}\tilde{y}(t)\big{\rrbracket},\ell),$
$\big{\llbracket}\tilde{u}_{i}(t)\big{\rrbracket}$ $=\big{\llbracket}\tilde{x}_{n_{i},n_{i}}(t)\big{\rrbracket}+\big{\llbracket}\tilde{D}_{i}\big{\rrbracket}\big{\llbracket}\tilde{y}(t)\big{\rrbracket},$
where
$\big{\llbracket}\tilde{A}_{n_{i}}\big{\rrbracket}\!=\!\begin{bmatrix}0&\cdots&0&\big{\llbracket}\tilde{a}_{n_{i},0}\big{\rrbracket}\\[1.0pt] 1&\cdots&0&\big{\llbracket}\tilde{a}_{n_{i},1}\big{\rrbracket}\\ \vdots&\ddots&\vdots&\vdots\\ 0&\cdots&1&\big{\llbracket}\tilde{a}_{n_{i},n_{i}-1}\big{\rrbracket}\end{bmatrix},\ \big{\llbracket}\tilde{x}_{n_{i}}(t)\big{\rrbracket}\!=\!\begin{bmatrix}\big{\llbracket}\tilde{x}_{n_{i},1}(t)\big{\rrbracket}\\[1.0pt] \big{\llbracket}\tilde{x}_{n_{i},2}(t)\big{\rrbracket}\\ \vdots\\ \big{\llbracket}\tilde{x}_{n_{i},n_{i}}(t)\big{\rrbracket}\end{bmatrix}$,
and $(a_{n_{i},0},\dots,\allowbreak a_{n_{i},n_{i}-1})$ are the coefficients of the characteristic polynomial of $A_{n_{i}}$. The computation of each subsystem requires $n_{i}+n_{i}p+p$ multiplications of shares, and hence, $\sum_{i=1}^{m}n_{i}+n_{i}p+p=(n+m)p+n$ Beaver triples are consumed to evaluate the $m$ subsystems. The total communication cost is then given as $2(3(n+m)p+5n+m+p)\ell_{q}$ bit, thereby reducing the communication complexity from the client to the servers to $O((n+m)p\log q)$. The modification also improves the computational complexities of the client and servers to $O((n+m)p)$.
6.2 Pseudorandom function
By definition, the shares for the party $P_{1}$ are random numbers in $\mathbb{Z}_{q}$ and independent of messages. This characteristic allows us to utilize a pseudorandom function further to reduce the communication cost between the client and parties. A pseudorandom function is a polynomial-time deterministic algorithm $\mathsf{PRF}:\{0,1\}^{\ell_{\mathrm{key}}(\lambda)}\times\{0,1\}^{\ell_{\mathrm{in}}(\lambda)}\to\{0,1\}^{\ell_{\mathrm{out}}(\lambda)}$, which takes an $\ell_{\mathrm{key}}(\lambda)$-bit key and an $\ell_{\mathrm{in}}(\lambda)$-bit string as input and outputs an $\ell_{\mathrm{out}}(\lambda)$-bit string, where $\lambda$ is a security parameter [28]. It is necessary that any polynomial-time algorithm without knowledge of the key cannot distinguish the output from a uniformly random string over $\{0,1\}^{\ell_{\mathrm{out}}(\lambda)}$.
Let $\mathsf{PRF}_{K}$ be a pseudorandom function with a key $K\in\{0,1\}^{\ell_{\mathrm{key}}(\lambda)}$ whose input and output domains are $\{0,1\}^{\ell_{\mathrm{in}}(\lambda)}$ and $\mathbb{Z}_{q}$, respectively. To generate shares for $P_{1}$, in the offline phase, the client randomly chooses a key $K$ and transmits $\mathsf{PRF}_{K}$ to $P_{1}$. In the online phase, the client and $P_{1}$ can generate the same pseudorandom number $r\leftarrow\mathsf{PRF}_{K}(t\bmod 2^{\ell_{\mathrm{in}}(\lambda)})$ regardless of $m\in\mathbb{Z}_{q}$. Then, $P_{1}$ sets $\llbracket m\rrbracket_{1}=r$, and the client computes the share for $P_{2}$ as $\llbracket m\rrbracket_{2}=m-r\bmod q$, which satisfies the requirement (P1). By applying this technique to each share generation, all communications from the client to $P_{1}$ can be eliminated. This reduces the total communication cost between the client and parties to $(3(n+m)p+5n+2m+p)\ell_{q}$ bit. Note that, because $\mathsf{PRF}_{K}$ is deterministic, the client must refresh the key $K$ every $\tau\in\mathbb{N}$ time step to ensure the security of $\mathsf{PRF}_{K}$. Although this process incurs additional communication costs, the modified share generation remains more efficient than the original one.
6.3 Pseudorandom correlation function
Instead of applying the pseudorandom functions, one can use a pseudorandom correlation function [36], which is analogous to a pseudorandom function that generates correlated randomness. A pseudorandom correlation function $\mathsf{PCF}$ is a two-party protocol in which each party $P_{i}$, $i\in\{1,2\}$ is given a key $K_{i}$ and, for some input string $x$, the parties can locally compute correlated strings $(\mathsf{PCF}_{K_{1}}(x),\mathsf{PCF}_{K_{2}}(x))$, where the outputs and random samples from the target correlation are indistinguishable [37].
In our scenario, the keys can be generated by the client, allowing the parties to obtain Beaver triples $(\llbracket a\rrbracket_{i},\llbracket b\rrbracket_{i},\llbracket c\rrbracket_{i})$ on demand during the online phase by invoking a pseudorandom correlation function $\mathsf{PCF}_{K_{i}}$ for the Beaver triple correlation with a common reference string $\mathsf{crs}$ (e.g., time $t$), namely $(\llbracket a\rrbracket_{i},\llbracket b\rrbracket_{i},\llbracket c\rrbracket_{i})\leftarrow\mathsf{PCF}_{K_{i}}(\mathsf{crs})$. Moreover, the shares of random numbers $(\llbracket r_{h}\rrbracket,\llbracket r^{\prime}_{h}\rrbracket)$ can be generated in the same manner. This approach eliminates the transmission of the auxiliary inputs $\mathsf{aux}_{i}$, thereby reducing the total communication cost between the client and parties to $2(m+p)\ell_{q}$ bit. It also decreases the computational complexity of the client to $O(m+p)$, as the client only needs to execute the $\mathsf{Share}$ and $\mathsf{Reconst}$ algorithms $p$ and $m$ times, respectively.
Integrating the proposed protocol with pseudorandom correlation functions is the most promising approach to significantly improve the communication complexity. In this pursuit, exploring efficient constructions of such pseudorandom correlation functions in real-time control systems is future work.
7 Numerical Examples
Consider a PID controller of parallel form in [14]. The controller parameters of (2) are $A=\quantity[\begin{smallmatrix}2-N_{d}&N_{d}-1\\ 1&0\end{smallmatrix}]$, $B=[1\ 0]^{\top}$, $C=[c_{1}\ c_{2}]$, and $D=d$, where $c_{1}=K_{i}T_{s}-N_{d}^{2}K_{d}/T_{s}$, $c_{2}=N_{d}K_{i}T_{s}-K_{i}T_{s}+N_{d}^{2}K_{d}/T_{s}$, and $d=K_{p}+N_{d}K_{d}/T_{s}$. Here, $K_{p}$, $K_{i}$, and $K_{d}$ are proportional, integral, and derivative gains, respectively, $N_{d}\in\mathbb{N}$ is the parameter of a derivative filter, and $T_{s}\in\mathbb{R}$ is a sampling time. Note that, if $2-N_{d},N_{d}-1\in\mathbb{Z}_{\langle k-\ell\rangle}$, the parameters $A$ and $B$ are already integers. Then, the proposed controller (12) can be simplified as
$\big{\llbracket}\tilde{x}(t+1)\big{\rrbracket}$ $=\big{\llbracket}A\big{\rrbracket}\big{\llbracket}\tilde{x}(t)\big{\rrbracket}+\big{\llbracket}B\big{\rrbracket}\big{\llbracket}\tilde{y}(t)\big{\rrbracket},$
$\big{\llbracket}\tilde{u}(t)\big{\rrbracket}$ $=\big{\llbracket}\tilde{C}\big{\rrbracket}\big{\llbracket}\tilde{x}(t)\big{\rrbracket}+\big{\llbracket}\tilde{D}\big{\rrbracket}\big{\llbracket}\tilde{y}(t)\big{\rrbracket},$
and the $\mathsf{Trunc}$ protocol is not necessary to implement (2).
Suppose that the plant (1) and sampling time are given by the single-input and single-output system in  [38, Eq. (2), with $\alpha=0.2$] and $T_{s}=0.1$ s, respectively. Let $\epsilon=2^{-10}$, $\lambda=80$ bit, $k-\ell=8$ bit, $x_{p,0}=[100\ 100\ 100\ 100]^{\top}$, and $x_{0}=[0\ 0]^{\top}$. For instance, the PID controller with the parameters $A=\quantity[\begin{smallmatrix}1&0\\ 1&0\end{smallmatrix}]$, $B=[1\ 0]^{\top}$, C=[2.73689272.96540833]C=[2.7368927\ {-2.96540833}], and D=5.01071167D=-5.01071167 satisfies Assumption 1, Assumption 2, and (15) with $\ell=32,40,48,56$ bit. The bit length of modulus $q$ can be chosen as $256$ bit to satisfy (14) for all the choices of $\ell$. Fig. 3 shows the input errors $\norm{u(t)-\hat{u}(t)}$. In this figure, the input errors remain below $\epsilon=2^{-10}$ for all $\ell=32,40,58,64$ and $t=0,\dots,50$, thereby supporting Theorem 5.5. Furthermore, the figure suggests that increasing the bit length $\ell$ can reduce the errors.
Figure 3: Input errors between the original and two-party PID controls with $\ell=32,40,48,56$ bit.
Next, we consider an observer-based controller for the plant (1) of multi-input and multi-output. The controller parameters of (2) are given as $A=A_{p}+B_{p}F-LC_{p}$, $B=L$, $C=F$, $D=\mathbf{0}_{m\times p}$, where $F$ and $L$ are feedback gains. Because the controller parameters are not necessarily integers, we used the proposed controller (12) with the $\mathsf{Trunc}$ protocol.
Suppose the plant is given as the four-tank process in [39] with $h_{1}^{0}=12.4$ cm, $h_{2}^{0}=12.7$ cm, $h_{3}^{0}=1.8$ cm, $h_{4}^{0}=1.4$ cm, $v_{1}^{0}=3$ V, $v_{2}^{0}=3$ V, $\gamma_{1}=0.7$, $\gamma_{2}=0.6$ and $T_{s}=0.5$ s. Let $\epsilon$, $\lambda$, $k$-$\ell$, and $\ell$ be as in the PID control example. Let $x_{p,0}=[10\ 10\ 10\ 10]^{\top}$ and $x_{0}=[0\ 0\ 0\ 0]^{\top}$. The controller with the parameters,
$A$ $=\begin{bmatrix}[r]\scriptstyle 0.56817627&\scriptstyle-0.00167847&\scriptstyle 0.01213074&\scriptstyle-0.00909424\\[-3.0pt] \scriptstyle-0.00201416&\scriptstyle 0.57826233&\scriptstyle-0.00939941&\scriptstyle 0.00976562\\[-3.0pt] \scriptstyle-0.15261841&\scriptstyle-0.01811218&\scriptstyle 0.97219849&\scriptstyle-0.00508118\\[-3.0pt] \scriptstyle-0.01197815&\scriptstyle-0.15417480&\scriptstyle-0.00314331&\scriptstyle 0.98011780\end{bmatrix}$,
$B$ $=\begin{bmatrix}[c]\scriptstyle 0.78367615&\scriptstyle 0&\scriptstyle 0.30230713&\scriptstyle 0\\[-3.0pt] \scriptstyle 0&\scriptstyle 0.78463745&\scriptstyle 0&\scriptstyle 0.30718994\end{bmatrix}^{\top}$,
$C$ $=\begin{bmatrix}[r]\scriptstyle-0.77249146&\scriptstyle-0.03674316&\scriptstyle-0.20259094&\scriptstyle-0.21775818\\[-3.0pt] \scriptstyle-0.06149292&\scriptstyle-0.76373291&\scriptstyle-0.29937744&\scriptstyle-0.21432495\end{bmatrix}$,
and a $256$-bit prime $q$ satisfy Assumption 1, Assumption 2, (14), and (16). Similar to the PID control example, Fig. 4 shows the input errors for $\ell=32,40,48,56$ bit. This figure demonstrates that the performance degradation in the proposed protocol is negligible even when using the truncation protocol.
Figure 4: Input errors between the original and two-party observer-based controls with $\ell=32,40,48,56$ bit.
8 Conclusion
We proposed a two-party protocol for the computation of a dynamic controller. The proposed protocol was realized based on a secret sharing scheme without the need for controller state decryption or input re-encryption and achieved statistical security. The performance of the dynamic controller under the proposed protocol can be made arbitrarily close to that of the original controller by approximating the controller parameters and initial state as sufficiently precise fixed-point numbers. We compared the computational and communication complexities of the protocol with those of conventional encrypted controllers and presented some improvements to the protocol using coordinate transformation and cryptographic primitives. Furthermore, the feasibility of the protocol was demonstrated through numerical examples.
In future work, we will extend the protocol to more advanced control strategies, such as model predictive control. We will also consider guaranteeing security not only against semi-honest adversaries but also against active adversaries who deviate from our intended behavior. This would address the limitation imposed by Assumption 3.
.1 Proof of Lemma 1
By the conditions of the bit lengths of $m$, $r$, and $r^{\prime}$, it follows from (P2) to (P4) that $\llbracket m_{r}\rrbracket=\llbracket m+2^{\ell}r+r^{\prime}+2^{\ell-1}\bmod q\rrbracket=\llbracket m+2^{\ell}r+r^{\prime}+2^{\ell-1}\rrbracket$. Thus, from (P1), we have $m_{r}=\mathsf{Reconst}(\llbracket m_{r}\rrbracket)=m+2^{\ell}r+r^{\prime}+2^{\ell-1}$. It then holds that $m_{r}-2^{\ell-1}\bmod 2^{\ell}=m+r^{\prime}\bmod 2^{\ell}=(m\bmod 2^{\ell})+(r^{\prime}\bmod 2^{\ell})+2^{\ell}w=(m\bmod 2^{\ell})+r^{\prime}+2^{\ell}w$, where $r^{\prime}\in\mathbb{Z}_{\langle\ell\rangle}$ and $w\in\{-1,0,1\}$. The properties from (P2) to (P4) yield
$\llbracket m^{\prime}\rrbracket$ $=\llbracket\operatorname{inv}(2^{\ell},q)(m+r^{\prime}-(m_{r}-2^{\ell-1}\bmod 2^{\ell}))\bmod q\rrbracket,$
$=\llbracket\operatorname{inv}(2^{\ell},q)(m-(m\bmod 2^{\ell})+2^{\ell}w)\bmod q\rrbracket,$
$=\left\llbracket\left\lceil\frac{m}{2^{\ell}}\right\rfloor+w\bmod q\right\rrbracket=\left\llbracket\left\lceil\frac{m}{2^{\ell}}\right\rfloor+w\right\rrbracket,$
[…]
.2 Proof of Lemma 5.3
We first note that $c$ and $\gamma$ in the claim exist from Assumption 2. In addition, $\alpha>1$, $\beta>1$, and $c/(1-\gamma)>1$ are fulfilled. Let $\kappa=\lfloor\log_{2}q\rfloor-\lambda-1$, and then, $\mathbb{Z}_{\langle k\rangle}\subset\mathbb{Z}_{\langle\kappa\rangle}\subset\mathbb{Z}_{q}$ holds because $\log_{2}q>\kappa+\lambda+1$ and $\kappa\geq k+1+\lfloor\log_{2}\max\{n,p\}\alpha\beta\frac{c}{1-\gamma}\rfloor$. This implies that $\tilde{A}=\bar{A}$, $\tilde{B}=\bar{B}$, $\tilde{C}=\bar{C}$, $\tilde{D}=\bar{D}$, and $\tilde{x}_{0}=\bar{x}_{0}$. From (P1), (P4), and (P5), if $\tilde{A}\tilde{x}(t)+\tilde{B}\tilde{y}(t)\in\mathbb{Z}_{\langle\kappa\rangle}^{n}$ and $\bar{y}(t)\in\mathbb{Z}_{q}^{p}$, Lemma 1 guarantees that
$\tilde{x}(t+1)$ $=\mathsf{Reconst}\quantity(\mathsf{Trunc}\quantity(\big{\llbracket}\tilde{A}\big{\rrbracket}\big{\llbracket}\tilde{x}(t)\big{\rrbracket}+\big{\llbracket}\tilde{B}\big{\rrbracket}\big{\llbracket}\tilde{y}(t)\big{\rrbracket},\ell)),$
$=\left\lceil\quantity(\bar{A}\tilde{x}(t)+\bar{B}\bar{y}(t))/2^{\ell}\right\rfloor+w(t).$
Similary, if $\tilde{C}\tilde{x}(t)+\tilde{D}\tilde{y}(t)\in\mathbb{Z}_{q}^{m}$ and $\bar{y}(t)\in\mathbb{Z}_{q}^{p}$, it follows that
$\tilde{u}(t)$ $=\mathsf{Reconst}\quantity(\big{\llbracket}\tilde{C}\big{\rrbracket}\big{\llbracket}\tilde{x}(t)\big{\rrbracket}+\big{\llbracket}\tilde{C}\big{\rrbracket}\big{\llbracket}\tilde{y}(t)\big{\rrbracket}),$
$=\bar{C}\tilde{x}(t)+\bar{D}\bar{y}(t).$
Hence, the claim holds if $\tilde{A}\tilde{x}(t)+\tilde{B}\tilde{y}(t)\in\mathbb{Z}_{\langle\kappa\rangle}^{n}$, $\tilde{C}\tilde{x}(t)+\tilde{D}\tilde{y}(t)\in\mathbb{Z}_{q}^{m}$, and $\bar{y}(t)\in\mathbb{Z}_{q}^{p}$ for all $t\in\mathbb{N}_{0}$. The remainder of this proof shows this through induction.
Let $\hat{x}(t)=2^{-\ell}\tilde{x}(t)$ and $\hat{\chi}(t)=[\hat{x}_{p}(t)^{\top}\ \hat{x}(t)^{\top}]^{\top}$. For $t=0$, it holds that
$\norm\big{\bar{y}(0)}_{\infty}\leq\norm\big{2^{\ell}\hat{y}(0)}_{\infty}+\frac{1}{2}\leq 2^{\ell}\norm\big{C_{p}}_{\infty}\norm\big{\hat{\chi}(0)}_{\infty}+\frac{1}{2}<\alpha\beta,$
and
$\max\quantity{\norm\big{\tilde{A}\tilde{x}(0)+\tilde{B}\tilde{y}(0)}_{\infty},\norm\big{\tilde{C}\tilde{x}(0)+\tilde{D}\tilde{y}(0)}_{\infty}}$
$\leq 2^{k-1}\max\{n,p\}\quantity(\norm*{\bar{x}_{0}}_{\infty}+\norm*{\bar{y}(0)}_{\infty}),$
$\leq 2^{k-1}\max\{n,p\}\quantity(2^{\ell}\quantity(1+\norm*{C_{p}}_{\infty})\norm*{\hat{\chi}(0)}_{\infty}+\frac{1}{2}),$
$<2^{k-1}\max\{n,p\}\alpha\beta,$
which means $\tilde{A}\tilde{x}(0)+\tilde{B}\tilde{y}(0)\in\mathbb{Z}_{\langle\kappa\rangle}^{n}$, $\tilde{C}\tilde{x}(0)+\tilde{D}\tilde{y}(0)\in\mathbb{Z}_{\langle\kappa\rangle}^{m}\subset\mathbb{Z}_{q}^{m}$, and $\tilde{y}(0)=\bar{y}(0)\in\mathbb{Z}_{\langle\kappa\rangle}^{p}\subset\mathbb{Z}_{q}^{p}$ because $\log_{2}\alpha\beta<k-1+\log_{2}\max\{n,p\}\alpha\beta<\kappa$.
Assume that $\tilde{A}\tilde{x}(t)+\tilde{B}\tilde{y}(t)\in\mathbb{Z}_{\langle\kappa\rangle}^{n}$, $\tilde{C}\tilde{x}(t)+\tilde{D}\tilde{y}(t)\in\mathbb{Z}_{q}^{m}$, and $\tilde{y}(t)=\bar{y}(t)\in\mathbb{Z}_{q}^{p}$ hold for all $t=0,\dots,\tau-1$ and $\tau\in\mathbb{N}$. With (P1), (P4), and (P5), it follows from (7) to (13) that, for all $t=0,\dots,\tau-1$,
$\hat{x}_{p}(t+1)$ $=A_{p}\hat{x}_{p}(t)+2^{-2\ell}B_{p}\tilde{u}(t),$
$=(A_{p}+B_{p}DC_{p})\hat{x}_{p}(t)+B_{p}C\hat{x}(t)+B_{p}Dd_{1}(t),$
$\hat{x}(t+1)$ $=2^{-\ell}\tilde{x}(t+1),$
$=BC_{p}\hat{x}_{p}(t)+A\hat{x}(t)+Bd_{1}(t)+d_{2}(t),$
where $d_{1}(t)=2^{-\ell}\bar{y}(t)-\hat{y}(t)$, $d_{2}(t)=2^{-\ell}(e(t)+w(t))$, and $e(t)=\lceil A\tilde{x}(t)+B\bar{y}(t)\rfloor-(A\tilde{x}(t)+B\bar{y}(t))$. Hence, we have
$\hat{\chi}(t+1)=\Phi\hat{\chi}(t)+\Gamma d_{1}(t)+\Xi d_{2}(t)$ (17)
for all $t=0,\dots,\tau-1$, where $\Xi=[\mathbf{0}_{n_{p}\times n}^{\top}\ I_{n}]^{\top}$. This implies $\hat{\chi}(\tau)=\Phi^{\tau}\hat{\chi}(0)+\sum_{s=0}^{\tau-1}\Phi^{\tau-1-s}\quantity(\Gamma d_{1}(s)+\Xi d_{2}(s))$.
Using this equation, the norm of $\tilde{x}(\tau)$ is bounded by
$\norm*{\tilde{x}(\tau)}_{\infty}=2^{\ell}\norm*{\hat{x}(\tau)}_{\infty}\leq 2^{\ell}\norm*{\hat{\chi}(\tau)}_{\infty},$
$\leq 2^{\ell}\quantity(\norm*{\hat{\chi}(0)}_{\infty}+\frac{2^{-\ell}}{2}\norm*{\Gamma}_{\infty}+\frac{2^{-\ell}\cdot 3}{2}\norm*{\Xi}_{\infty})\sum_{s=0}^{\tau}\norm*{\Phi^{s}}_{\infty},$
$<\quantity(2^{\ell}\norm*{\hat{\chi}(0)}_{\infty}+\frac{1}{2}\norm*{\Gamma}_{\infty}+\frac{3}{2})\frac{c}{1-\gamma}=\beta\frac{c}{1-\gamma},$
where $\norm{d_{1}(s)}_{\infty}\leq 2^{-\ell}\cdot(1/2)$, $\norm{d_{2}(s)}_{\infty}\leq 2^{-\ell}\cdot(3/2)$, and $\sum_{s=0}^{k}\norm{\Phi^{s}}_{\infty}\leq\sum_{s=0}^{k}\norm{\Phi^{s}}\leq\sum_{s=0}^{k}c\gamma^{s}<\sum_{s=0}^{\infty}c\gamma^{s}=c/(1-\gamma)$. Moreover, the norm of $\bar{y}(\tau)$ is bounded by
$\norm*{\bar{y}(\tau)}_{\infty}$ $\leq 2^{\ell}\norm*{C_{p}}_{\infty}\norm*{\hat{\chi}(\tau)}_{\infty}+\frac{1}{2}\leq\norm*{C_{p}}_{\infty}\beta\frac{c}{1-\gamma}+\frac{1}{2},$
$<\quantity(\norm*{C_{p}}_{\infty}+\frac{1}{2})\beta\frac{c}{1-\gamma}<\alpha\beta\frac{c}{1-\gamma},$
which means $\tilde{y}(\tau)=\bar{y}(\tau)\in\mathbb{Z}_{\langle\kappa\rangle}^{p}\subset\mathbb{Z}_{q}^{p}$ because
$\alpha\beta\frac{c}{1-\gamma}$ $<2^{k-1}\max\{n,p\}\alpha\beta\frac{c}{1-\gamma},$
$<2^{k-1}\quantity(\left\lfloor\max\{n,p\}\alpha\beta\frac{c}{1-\gamma}\right\rfloor+1),$
$\leq 2^{k-1+\lfloor\max\{n,p\}\alpha\beta c/(1-\gamma)\rfloor}=2^{\kappa-2}<2^{\kappa-1}-1.$
From these bounds, it holds that
$\max\quantity{\norm\big{\tilde{A}\tilde{x}(\tau)+\tilde{B}\tilde{y}(\tau)}_{\infty},\norm\big{\tilde{C}\tilde{x}(\tau)+\tilde{D}\tilde{y}(\tau)}_{\infty}}$
$\leq 2^{k-1}\max\{n,p\}\quantity(\norm{\tilde{x}(\tau)}_{\infty}+\norm{\bar{y}(\tau)}_{\infty}),$
$<2^{k-1}\max\{n,p\}\quantity(1+\norm*{C_{p}}_{\infty}+\frac{1}{2})\beta\frac{c}{1-\gamma},$
$=2^{k-1}\max\{n,p\}\alpha\beta\frac{c}{1-\gamma}<2^{\kappa-1}-1.$
Therefore, $\tilde{A}\tilde{x}(\tau)+\tilde{B}\tilde{y}(\tau)\in\mathbb{Z}_{\langle\kappa\rangle}^{n}$ and $\tilde{C}\tilde{x}(\tau)+\tilde{D}\tilde{y}(\tau)\in\mathbb{Z}_{\langle\kappa\rangle}^{m}\subset\mathbb{Z}_{q}^{m}$ are satisfied.
.3 Proof of Theorem 5.5
Let $\hat{x}(t)=2^{\ell}\tilde{x}(t)$, $\chi(t)=[x_{p}(t)^{\top}\ x(t)^{\top}]^{\top}$, and $\hat{\chi}(t)=[\hat{x}_{p}(t)^{\top}\ \hat{x}(t)^{\top}]^{\top}$. Define $\varepsilon(t)=\chi(t)-\hat{\chi}(t)$. If (14) is satisfied, from (3) and (17), the dynamics of $\varepsilon$ is given as
$\varepsilon(t+1)=\Phi\varepsilon(t)-\Gamma d_{1}(t)-\Xi d_{2}(t),\quad\varepsilon(0)=\mathbf{0}_{(n_{p}+n)\times 1},$
where $\Xi$, $d_{1}(t)$, and $d_{2}(t)$ are the same as in the proof of Lemma 5.3. From Assumption 2 and the definitions of $d_{1}(t)\in\mathbb{R}^{p}$ and $d_{2}(t)\in\mathbb{R}^{n}$, it holds that, for all $t\in\mathbb{N}_{0}$, $\norm{\Phi^{t}}\leq c\gamma^{t}$, $\norm{d_{1}(t)}\leq 2^{-\ell-1}\sqrt{p}$, and $\norm{d_{2}(t)}<2^{-\ell+1}\sqrt{n}$. Applying these bounds to the norm of $\varepsilon(t)$, we obtain
$\norm{\varepsilon(t)}$ $=\norm{\Phi^{t}\varepsilon(0)+\sum_{s=0}^{t-1}\Phi^{t-1-s}\quantity(\Gamma d_{1}(s)+\Xi d_{2}(s))},$
$\leq 2^{-\ell}\quantity(\frac{\sqrt{p}}{2}\norm*{\Gamma}+2\sqrt{n}\norm*{\Xi})\sum_{s=0}^{t-1}\norm*{\Phi^{s}},$
$<2^{-\ell}\quantity(\frac{\sqrt{p}}{2}\norm{\Gamma}+2\sqrt{n})\frac{c}{1-\gamma},$
where $\sum_{s=0}^{t-1}\norm*{\Phi^{s}}<c/(1-\gamma)$.
From Lemma 5.3, (2), and (13), the error of the control inputs is given as
$u(t)-\hat{u}(t)$ $=(Cx(t)+Dy(t))-2^{-2\ell}\quantity(\bar{C}\tilde{x}(t)+\bar{D}\bar{y}(t)),$
$=DC_{p}e_{p}(t)+Ce(t)+Dd_{1}(t),$
$=\Upsilon\varepsilon(t)+Dd_{1}(t).$
With the bounds of $\varepsilon(t)$ and $d_{1}(t)$, it concludes that
$\norm*{u(t)-\hat{u}(t)}\leq\norm{\Upsilon}\norm{\varepsilon(t)}+\norm{D}\norm{d_{1}(t)},$
$\leq 2^{-\ell}\quantity(\frac{\sqrt{p}}{2}\norm{\Gamma}\norm{\Upsilon}+2\sqrt{n}\norm{\Upsilon})\frac{c}{1-\gamma}+2^{-\ell-1}\sqrt{p}\norm{D},$
$<2^{-\ell}\quantity[\frac{\sqrt{p}}{2}\quantity(\norm{\Gamma}\norm{\Upsilon}+\norm{D})+2\sqrt{n}\norm{\Upsilon}]\frac{c}{1-\gamma}.$
Therefore, (4) holds if $\ell$ satisfies (15).
.4 Proof of
The claim can be obtained by applying the bounds $\norm{\Gamma}=\norm{B}\leq 2^{k-\ell-1}\sqrt{np}$, $\norm{\Upsilon}=\norm{C}\leq 2^{k-\ell-1}\sqrt{mn}$, and $\norm{D}\leq 2^{k-\ell-1}\sqrt{mp}$ to (15).